OpenStudy (anonymous):

does anyone wanna help me with my algebra work? ill give a testimonial and medal!(:

OpenStudy (anonymous):

What is the problem?

OpenStudy (therealmeeeee):

sure show me

OpenStudy (anonymous):

Like the equation is what?

OpenStudy (anonymous):

one second let me get them.

OpenStudy (anonymous):

Okay

OpenStudy (therealmeeeee):

kk

OpenStudy (anonymous):

here is the first two

OpenStudy (anonymous):

@TheRealMeeeee

OpenStudy (anonymous):

@Aimee98

OpenStudy (therealmeeeee):

yes

OpenStudy (therealmeeeee):

idk how too do that one

OpenStudy (johnweldon1993):

Slope = rise/run :)

OpenStudy (anonymous):

well shattt let m get the others

OpenStudy (johnweldon1993):

Use that for the first one....and for the second one...use the expression \[\large \text {slope} = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (solomonzelman):

\[\frac{y_1-y_2}{x_1-x_2}=slope\]use this, just plug in your points.

OpenStudy (anonymous):

:p

OpenStudy (anonymous):

thank yall! (: ill give all that helped testimonials

OpenStudy (johnweldon1993):

varying directly just means the equation y = kx you are given a (x,y) so plug those in to solve for 'k'....then use the same equation y = kx with your newly found 'k'....and plug in your x = 10 and solve for y

OpenStudy (isaiah.feynman):

Have you even tried solving these?

OpenStudy (johnweldon1993):

slope intercept form is y = mx + b where 'm' = the slope and b = the y-intercept...so just plug in your 2 numbers there...

OpenStudy (johnweldon1993):

You need to find the slope given 2 points again (look above) then just plug one of your points into the rest of the equation y = mx + b

OpenStudy (anonymous):

yes, ive failed all the math ive taken my whole life. its almost impossibe for me to learn now. but john thank you !

OpenStudy (johnweldon1993):

Anytime :)

OpenStudy (isaiah.feynman):

Oh sorry.